\begin{thebibliography}{10}

\bibitem{ak-chou-jean-18}
{\sc Aksamit, A., Choulli, T., and Jeanblanc, M.}
\newblock Thin times and random times' decomposition.
\newblock {\em Electron. J. Probab. 26\/} (2021), Paper No. 31, 22.

\bibitem{aksamit-jean-17}
{\sc Aksamit, A., and Jeanblanc, M.}
\newblock {\em Enlargement of filtration with finance in view}.
\newblock SpringerBriefs in Quantitative Finance. Springer, Cham, 2017.

\bibitem{band-calvia-col22}
{\sc Bandini, E., Calvia, A., and Colaneri, K.}
\newblock Stochastic filtering of a pure jump process with predictable jumps
  and path-dependent local characteristics.
\newblock {\em Stochastic Process. Appl. 151\/} (2022), 396--435.

\bibitem{ba-con-ditella-21}
{\sc Bandini, E., Confortola, F., and Di~Tella, P.}
\newblock On the compensator of step processes in progressively enlarged
  filtrations and related control problems.
\newblock {\em ALEA Lat. Am. J. Probab. Math. Stat. 21}, 1 (2024).

\bibitem{MR2070167}
{\sc B\'{e}langer, A., Shreve, S.~E., and Wong, D.}
\newblock A general framework for pricing credit risk.
\newblock {\em Math. Finance 14}, 3 (2004), 317--350.

\bibitem{MR1406564}
{\sc Bertoin, J.}
\newblock {\em L\'{e}vy processes}, vol.~121 of {\em Cambridge Tracts in
  Mathematics}.
\newblock Cambridge University Press, Cambridge, 1996.

\bibitem{bla-jean04}
{\sc Blanchet-Scalliet, C., and Jeanblanc, M.}
\newblock Hazard rate for credit risk and hedging defaultable contingent
  claims.
\newblock {\em Finance Stoch. 8}, 1 (2004), 145--159.

\bibitem{Bre-yor78}
{\sc Br{\'e}maud, P., and Yor, M.}
\newblock Changes of filtrations and of probability measures.
\newblock {\em Z. Wahrsch. Verw. Gebiete 45}, 4 (1978), 269--295.

\bibitem{ca-jean-za13}
{\sc Callegaro, G., Jeanblanc, M., and Zargari, B.}
\newblock Carthaginian enlargement of filtrations.
\newblock {\em ESAIM Probab. Stat. 17\/} (2013), 550--566.

\bibitem{caltor-accessible-23}
{\sc Calzolari, A., and Torti, B.}
\newblock Martingale representation on enlarged filtrations: the role of the
  accessible jump times.
\newblock Tech. rep., arXiv:1708.05858v3 [math.PR], 2023.

\bibitem{jean-coku-nike12}
{\sc Coculescu, D., Jeanblanc, M., and Nikeghbali, A.}
\newblock Default times, no-arbitrage conditions and changes of probability
  measures.
\newblock {\em Finance Stoch. 16}, 3 (2012), 513--535.

\bibitem{davis2}
{\sc Davis, M. H.~A.}
\newblock Martingale representation and all that.
\newblock In {\em Advances in control, communication networks, and
  transportation systems}, Systems Control Found. Appl. Birkh\"auser Boston,
  Boston, MA, 2005, pp.~57--68.

\bibitem{davis1}
{\sc Davis, M. H.~A., and Varaiya, P.}
\newblock The multiplicity of an increasing family of {$\sigma $}-fields.
\newblock {\em Ann. Probability 2\/} (1974), 958--963.

\bibitem{del-me-a}
{\sc Dellacherie, C., and Meyer, P.-A.}
\newblock Probabilities and potential. {A}.
\newblock vol.~29 of {\em North-Holland Mathematics Studies}. North-Holland
  Publishing Co., Amsterdam, 1978, pp.~xvii+463.

\bibitem{ditella-eng-22}
{\sc Di~Tella, P., and Engelbert, H.-J.}
\newblock Martingale representation in progressively enlarged {L}\'{e}vy
  filtrations.
\newblock {\em Stochastics 94}, 2 (2022), 311--333.

\bibitem{elka-jean-jiao09}
{\sc El~Karoui, N., Jeanblanc, M., and Jiao, Y.}
\newblock What happens after a default: the conditional density approach.
\newblock {\em Stochastic Process. Appl. 120}, 7 (2010), 1011--1032.

\bibitem{fontana-smith17}
{\sc Fontana, C., and Schmidt, T.}
\newblock General dynamic term structures under default risk.
\newblock {\em Stochastic Process. Appl. 128}, 10 (2018), 3353--3386.

\bibitem{MR3758922}
{\sc Gehmlich, F., and Schmidt, T.}
\newblock Dynamic defaultable term structure modeling beyond the intensity
  paradigm.
\newblock {\em Math. Finance 28}, 1 (2018), 211--239.

\bibitem{MR4529908}
{\sc Gueye, D., and Jeanblanc, M.}
\newblock Generalized {C}ox model for default times.
\newblock {\em Front. Math. Finance 1}, 4 (2022), 467--489.

\bibitem{he-wang-yan92}
{\sc He, S.~W., Wang, J.~G., and Yan, J.~A.}
\newblock Semimartingale theory and stochastic calculus.
\newblock Kexue Chubanshe (Science Press), Beijing; CRC Press, Boca Raton, FL,
  1992, pp.~xiv+546.

\bibitem{jacod}
{\sc Jacod, J.}
\newblock Calcul stochastique et probl\`emes de martingales.
\newblock vol.~714 of {\em Lecture Notes in Mathematics}. Springer, Berlin,
  1979, pp.~x+539.

\bibitem{jacod85}
{\sc Jacod, J.}
\newblock Grossissement initial, hypoth\`ese ({H}{'}), et th\'eor\`eme de
  {G}irsanov.
\newblock In {\em {T}hierry {J}eulin and { M}arc {Y}or editors, Grossissements
  de Filtrations: Exemples et Applications}, vol.~1118 of {\em Lecture Notes in
  Math.} Springer, Berlin, 1985, pp.~15--35.

\bibitem{jiao-li15}
{\sc Jiao, Y., and Li, S.}
\newblock The generalized density approach in progressive enlargement of
  filtrations.
\newblock {\em Electron. J. Probab. 20\/} (2015), no. 85, 21.

\bibitem{MR3758923}
{\sc Jiao, Y., and Li, S.}
\newblock Modeling sovereign risks: from a hybrid model to the generalized
  density approach.
\newblock {\em Math. Finance 28}, 1 (2018), 240--267.

\bibitem{MR4152276}
{\sc Kardaras, C., and Ruf, J.}
\newblock Filtration shrinkage, the structure of deflators, and failure of
  market completeness.
\newblock {\em Finance Stoch. 24}, 4 (2020), 871--901.

\bibitem{kusuoka99}
{\sc Kusuoka, S.}
\newblock A remark on default risk models.
\newblock In {\em Advances in mathematical economics, {V}ol. 1 ({T}okyo,
  1997)}, vol.~1 of {\em Adv. Math. Econ.} Springer, Tokyo, 1999, pp.~69--82.

\bibitem{Prott}
{\sc Protter, P.~E.}
\newblock Stochastic integration and differential equations.
\newblock Second~ed., vol.~21 of {\em Applications of Mathematics (New York)}.
  Springer-Verlag, Berlin, 2004, pp.~xiv+415.
\newblock Stochastic Modelling and Applied Probability.

\end{thebibliography}

\documentclass[dvips, 12pt, a4paper]{paper}
\usepackage{amsmath,amssymb,amsthm, dsfont}
\usepackage{color}
\definecolor{violet}{rgb}{0.5,0,0.8}

\theoremstyle{plain}
\newtheorem{thm}{Theorem}[section]

\theoremstyle{plain}
\newtheorem{cor}[thm]{Corollary}

\theoremstyle{plain}
\newtheorem{lemma}[thm]{Lemma}

\theoremstyle{plain}

\theoremstyle{plain}
\newtheorem{proposition}[thm]{Proposition}

\theoremstyle{plain}
\newtheorem{definition}[thm]{Definition}

\theoremstyle{plain}

\theoremstyle{plain}
\newtheorem{remark}[thm]{Remark}

\begin{document}
\title{Thin-thick approach to martingale representations on progressively enlarged filtrations}
\author{Antonella Calzolari \thanks{Dipartimento di Matematica - Universit\`a di Roma
``Tor Vergata'', via della Ricerca Scientifica 1, I 00133 Roma,
Italy }  \and Barbara Torti $^*$}
\maketitle
\begin{abstract}
We study the predictable representation property in the progressive enlargement $\mathbb{F}^\tau$ of a reference filtration $\mathbb{F}$ by a random time $\tau$.~Our approach is based on the decomposition of any random time into two parts, one overlapping $\mathbb{F}$-stopping times (thin part) and the other one that avoids $\mathbb{F}$-stopping times (thick part).~We assume that the $\mathbb{F}$-thin part of $\tau$ is nontrivial and prove a martingale representation theorem on $\mathbb{F}^\tau$.~We thus extend previous  results dealing with $\mathbb{F}$-avoiding random times.~We collect some examples of application to the enlargement of the natural filtration of a L\'evy process.
\end{abstract}
\keywords{Predictable representations property;
enlargement of filtration;
L\'evy processes;
compensators.}
\section{Introduction}\label{sec1}
Let $\mathbb{F}$ be a filtration and  $(M^j)_{j\in J}$, $J\subset\mathbb{N}$, a set of square-integrable pairwise orthogonal $\mathbb{F}$-martingales.~Assume that any
square-integrable $\mathbb{F}$-martingale can be represented as sum of stochastic integrals of predictable processes
w.r.t.~the elements of $(M^j)_{j\in J}$.~In other words $(M^j)_{j\in J}$ is an \textit{$\mathbb{F}$-basis} in the sense of Davis and Varaiya (see \cite{davis1}).~It follows that  the \textit{(strong) predictable representation property in $\mathbb{F}$} holds (see \cite{jacod}).\bigskip\\
Let now $\tau$ be a random time and let $\mathbb{F}^\tau$ be equal, up to standardization, to the filtration  $\mathbb{F}\vee\sigma(\tau\wedge\cdot)$.~$\mathbb{F}^\tau$ is known as the \textit{progressive enlargement of $\mathbb{F}$ by
$\tau$}.~Thus, $\tau$ is an $\mathbb{F}^\tau$ stopping time and it makes sense to consider the \textit{$\mathbb{F}^\tau$-compensated occurrence process of $\tau$}.~It deals with the $\mathbb{F}^\tau$-compensation in Doob's sense of the sub-martingale $\mathbb{I}_{\tau\leq }$.~In the following, we will denote it by $H^{\tau,\mathbb{F}^\tau}$.\bigskip\\
In this paper, we focus on the predictable representation property in $\mathbb{F}^\tau$.~More precisely, we study under which conditions the family of martingales $\left((M^j)_{j\in J},
H^{\tau,\mathbb{F}^\tau}\right)$ allows one to represent all $\mathbb{F}^\tau$-local martingales.\bigskip\\
Our hypotheses are less restrictive than those typically adopted in the papers dealing with this problem.~Most authors, in fact, assume that the
graph of $\tau$ is disjoint from the graph of any $\mathbb{F}$-stopping time or equivalently that the \textit{$\mathbb{F}$-avoiding condition} holds (see e.g.~\cite{kusuoka99}, \cite{bla-jean04},  \cite{elka-jean-jiao09}, \cite{jean-coku-nike12}, \cite{ca-jean-za13}, \cite{ditella-eng-22}).~Here instead, according to the \textit{thin-thick decomposition} of any random time (see \cite{ak-chou-jean-18}), $\tau$ may coincide with $\mathbb{F}$-stopping times.\bigskip\\
Therefore, we split the problem of the identification of an $\mathbb{F}^\tau$-basis into two steps:\begin{itemize}  \item [(i)] the construction of a basis on an ``intermediate filtration'' obtained progressively enlarging $\mathbb{F}$ by the \textit{thin part of $\tau$}; \item [(ii)] the construction of a basis on the progressive enlargement of the intermediate filtration by the \textit{thick part of $\tau$}.\end{itemize}
We stress that solving the first issue is the real goal of our work.\bigskip\\
Our main theorem reads as follows: $\left((M^j)_{j\in J},
H^{\tau,\mathbb{F}^\tau}\right)$ is an $\mathbb{F}^\tau$-basis as soon as the thin part of $\tau$ is composed by $\mathbb{F}$-predictable stopping times on which $\mathbb{F}$ is continuous, and any $\mathbb{F}$-martingale is also an  $\mathbb{F}^\tau$-martingale (see Theorem \ref{thm:new}).\bigskip\\
The theorem applies when $\mathbb{F}$ is the natural filtration of any L\'evy process since, in this case, $\mathbb{F}$ is a quasi-left continuous filtration (see Theorem \ref{cor:Levy}) and, in particular, when $\mathbb{F}$ is the natural filtration of a Brownian motion (see Corollary \ref{cor:Brow}).\bigskip\\
Results about the propagation of the martingale representation property to progressively enlarged filtrations are useful in many applied fields.~In credit risk theory
the random time $\tau$ plays the role of default time.~Financial models in which the default time may coincide with predictable stopping times have been considered first in \cite{MR2070167} and then  in \cite{MR3758923},  \cite{MR3758922} and \cite{fontana-smith17}.~Martingale representations theorems on extended filtrations may be crucial also when solving optimal stochastic control problems or filtering problems (see e.g.~\cite{ba-con-ditella-21} and \cite{band-calvia-col22}).\bigskip\\
%
%
%
This paper is organized as follows.~In Section 2 we fix the notations and recall some basic definitions.~In Section 3 we introduce the hypothesis on the thin part of $\tau$ and study, under this hypothesis, its compensated occurrence process (see Proposition \ref{lemma:eq:Delta}).~In Section 4, after showing the pairwise orthogonality of the set $\left((M^j)_{j\in J},
H^{\tau,\mathbb{F}^\tau}\right)$ (see Lemma \ref{cor}), we prove our main theorem (see Theorem \ref{thm:new}).~In Section 5, when  $\mathbb{F}$ is the natural filtration of a L\'evy process, we discuss some examples of application of the theorem.~Finally, we devote  Section 6 to brief comments and research proposals.
\section{Notations and definitions}\label{sec2}
Let  $(\Omega, \mathcal{F}, P)$ be a complete probability space and  $\mathbb{F}=(\mathcal{F}_t)_t$ a standard filtration on it.~$\mathcal{M}^2(P,\mathbb{F})$ denotes the Hilbert space of square-integrable $(P,\mathbb{F})$-martingales with inner product  $(M^j,M^l)\rightarrow E^P[M^j,M^l]_\infty$.\bigskip\\
Let us introduce the notion of the \textit{basis} of a filtration in the sense of Davis and Varaiya (see \cite{davis1}).\vspace{.1em}\\
 \begin{definition}\label{def:basis}
An $\mathbb{F}$-basis is a subset $(M^j)_{j\in J}, J\subset \mathbb{N},$ of $\mathcal{M}^2(P,\mathbb{F})$ whose elements are pairwise strongly orthogonal martingales
 such that each $V\in \mathcal{M}^2(P,\mathbb{F})$ satisfies
\begin{align*}
V_t=V_0+\sum_{j\in J}\int_0^t \Phi^V_j(s) d M^j_s,\;\;t\geq 0,
\end{align*}
where $V_0$ is a random variable $\mathcal{F}_0$-measurable and, for all $j\in J$,
$\Phi^V_j$ is a
predictable process that verifies $E^P\left[\int_0^\infty\,(\Phi^V_j)^2(s) d[M^j]_s\right]<+\infty$.
\end{definition}
\vspace{.1em}
\noindent We recall some basic facts about filtrations and random times.\bigskip\\
We call \textit{nontrivial} an $\overline{\mathbb{R}}^+$-valued random time $\tau$, such that $P(\tau< +\infty)>0$.\bigskip\\
Given a nontrivial random time $\tau$ the filtration $\sigma(\tau\wedge \cdot)$ is the minimal standard filtration which makes $\tau$ a stopping time.~We denote by
$\mathbb{F}^\tau$ the standard \textit{progressive enlargement of $\mathbb{F}$ by $\tau$}, that is
$$\mathcal{F}^\tau_t:=\bigcap_{s>t}\mathcal{F}_s\vee\sigma(\tau\wedge s).$$
As well-known, any $\mathbb{F}$-stopping time $\tau$ satisfies
 \begin{equation*}
 \tau=\tau^{a, \mathbb{F}}\wedge\tau^{i, \mathbb{F}}
 \end{equation*}
where $\tau^{a, \mathbb{F}}$ and $\tau^{i, \mathbb{F}}$ are the \textit{$\mathbb{F}$-accessible component of $\tau$} and the \textit{$\mathbb{F}$-totally inaccessible component of $\tau$}, respectively (see e.g.~Theorem 3,
 page 104, in \cite{Prott}).~More precisely,
there exist two disjoint events $A$ and $B$ such that  $P$-a.s.
$$A\cup B=(\tau<\infty)$$
 and
 \begin{equation}\label{eq-time-decomposition}
   \tau^{a, \mathbb{F}}=\tau\, \mathbb{I}_A + \infty\, \mathbb{I}_{A^c},\;\;\;\tau^{i, \mathbb{F}}=\tau\, \mathbb{I}_B + \infty\, \mathbb{I}_{B^c}.
 \end{equation}
Therefore,
 $$\tau\; \mathbb{I}_{\tau<+\infty}=\tau^{a, \mathbb{F}}\, \mathbb{I}_A + \tau^{i, \mathbb{F}}\, \mathbb{I}_B.$$
If $\sigma$ is any $\mathbb{F}$-predictable stopping time, then by definition
 \begin{equation*}
 P(\tau^{i, \mathbb{F}}=\sigma<+\infty)=0,
 \end{equation*}
As far as $\tau^{a, \mathbb{F}}$ is concerned,
 \begin{equation}\label{eq: accessible-r-t}
 [[\tau^{a, \mathbb{F}}]]\subset\bigcup_m [[\tau^{a, \mathbb{F}}_m]],
 \end{equation}
where $(\tau^{a, \mathbb{F}}_m)_{m}$ is a sequence of $\mathbb{F}$-predictable stopping times.~Such a sequence is not unique and it may be chosen in such a way that the corresponding graphs are pairwise disjoint (see Theorem 3.31, page 95, in \cite{he-wang-yan92}).~Any sequence
of $\mathbb{F}$-predictable stopping times  $(\tau^{a, \mathbb{F}}_m)_{m}$ satisfying (\ref{eq: accessible-r-t}) and with pairwise disjoint graphs takes the name of \textit{enveloping sequence of $\tau^{a, \mathbb{F}}$}.\vspace{.1em}\\
\begin{remark}
Note that if  $\mathbb{B}$ is a filtration different from $\mathbb{F}$ but such that $\tau$ is also $\mathbb{B}$-stopping time, then the $\mathbb{B}$-accessible component $\tau^{a, \mathbb{B}}$ is in general different from $\tau^{a, \mathbb{F}}$ and analogously $\tau^{i, \mathbb{B}}$ is in general different from $\tau^{i, \mathbb{F}}$.~Accessibility (totally inaccessibility) may get lost by restriction (expansion) of the filtration, that is an accessible (totally inaccessible) stopping time may be totally inaccessible (accessible)  w.r.t.~a smaller (bigger) filtration.
\end{remark}
\vspace{.1em}
\noindent In the next we refer to $\tau^{a,\sigma(\tau\wedge\cdot)}$ as the \textit{naturally accessible component of $\tau$}, and analogously to $\tau^{i,\sigma(\tau\wedge\cdot)}$ as the \textit{naturally totally inaccessible component of $\tau$}.\bigskip\\
A random time $\tau$ is \textit{$\mathbb{F}$-accessible} when $P(\tau=\tau^{a,\mathbb{F}})=1$, and analogously $\tau$ is  \textit{$\mathbb{F}$-totally inaccessible}, when $P(\tau=\tau^{i,\mathbb{F}})=1$ (if $\mathbb{F}=\sigma(\tau\wedge\cdot)$  \textit{naturally accessible} and \textit{naturally totally inaccessible}, respectively).\vspace{.1em}\\
\begin{remark}\label{rem:predictability}
It is worthwhile to recall that a naturally accessible stopping time cannot be naturally predictable (unless it coincides with a positive constant) and has an atomic law.~The law of a naturally totally inaccessible stopping time, instead, is diffusive (see Theorem IV-107, page 241, in \cite{del-me-a}).
\end{remark}
\vspace{.1em}
\noindent In the following, for any $\mathbb{F}$-stopping time $\tau$, $H^{\tau,\mathbb{F}}$ stays for  the \textit{$\mathbb{F}$-compensated occurrence process of $\tau$}, that is the $\mathbb{F}$-martingale obtained by $\mathbb{F}$-compensation of $\mathbb{I}_{\tau\leq \cdot}$~In formulas,
$$H^{\tau,\mathbb{F}}:=\mathbb{I}_{\tau\leq\cdot}-A^{\tau,\mathbb{F}},$$
where $A^{\tau,\mathbb{F}}$ is the $\mathbb{F}$-predictable compensator of $\mathbb{I}_{\tau\leq \cdot}$ (briefly, \textit{$\mathbb{F}$-compensator} of $\tau$).\vspace{.1em}\\
\begin{remark}\label{rem pred}
If $\tau$ is $\mathbb{F}$-predictable, then trivially $H^{\tau,\mathbb{F}}\equiv 0$.
\end{remark}
\vspace{.1em}
\begin{definition}\label{def:avoidance}
A random time $\tau$ satisfies Hypothesis ($\bold{\mathcal{A}}$) w.r.t.~$\mathbb{F}$ if $\tau$ avoids $\mathbb{F}$-stopping times, that is, if  $P(\tau = T < +\infty) = 0$ for every $\mathbb{F}$-stopping time $T$.
\end{definition}
\vspace{.1em}
\begin{remark}\label{rem:avoidance tot.in.ty}
If $\tau$ satisfies Hypothesis ($\bold{\mathcal{A}}$) w.r.t.~$\mathbb{F}$, then the $\mathbb{F}^\tau$-compensator of $\tau$ is continuous (see Lemma 3.6 in \cite{jean-coku-nike12}) and therefore $\tau$ is $\mathbb{F}^\tau$-totally inaccessible (see Theorem 1.43 in \cite{aksamit-jean-17}).
\end{remark}
\vspace{.1em}
\noindent A random time $\tau$ is called \textit{$\mathbb{F}$-thin} if
 $$\mathbb{I}_{\{\tau<\infty\}}\tau=\sum_{n}  T_n \mathbb{I}_{C_n},$$ where $T_n, n\geq 1,$ are  $\mathbb{F}$-stopping
times with disjoint graphs and, for all $n\geq 1$,
\begin{equation}\label{def:C_n} C_n=(\tau=T_n<\infty).\end{equation}
\noindent The sequence $(T_n)_{n\geq 1}$ is called an \textit{$\mathbb{F}$-exhausting sequence of $\tau$}, and  the event $C_n$ belongs to $\mathcal{F}^{\tau}_\infty$.\bigskip\\
For any random time $\tau$ the \textit{$\mathbb{F}$-thin-thick decomposition} holds, that is
 \begin{equation}\label{eq-eta-dec}
 \tau=\tau_1\wedge\tau_2,
 \end{equation}
where $\tau_1$ is $\mathbb{F}$-thin and is called the \textit{$\mathbb{F}$-thin part} of $\tau$, and $\tau_2$ satisfies Hypothesis ($\bold{\mathcal{A}}$) w.r.t.~$\mathbb{F}$ and is called the \textit{$\mathbb{F}$-thick part} of $\tau$ (see \cite{ak-chou-jean-18}).\\
Any $\mathbb{F}$-stopping time coincides with its $\mathbb{F}$-thin part and has a trivial $\mathbb{F}$-thick part.\vspace{.1em}\\
\begin{remark}\label{rem:composition}
$\mathbb{F}^{\tau}$ coincides with $\big(\mathbb{F}^{\tau_1}\big)^{\tau_2}=\big(\mathbb{F}^{\tau_2}\big)^{\tau_1}$  (see Theorem 5.1 in \cite{ak-chou-jean-18}).
\end{remark}
\vspace{.1em}
\noindent As usual, we write \textit{$\mathbb{F}\hookrightarrow\mathbb{F}^\tau$}, when any $\mathbb{F}$-local martingale is also a $\mathbb{F}^\tau$-local martingale, that is when the \textit{immersion property of  $\mathbb{F}$ in $\mathbb{F}^\tau$} holds.
\section{Some auxiliary results}
In this section, we introduce the class of those random times, that may overlap $\mathbb{F}$-predictable stopping times only.~When  $\mathbb{F}$ is the Brownian filtration all random times belong to this class.\vspace{.1em}\\
\begin{definition}\label{def:P}
A random time $\tau$ satisfies Hypothesis ($\bold{\mathcal{P}}$) w.r.t.~$\mathbb{F}$ when it admits a nontrivial $\mathbb{F}$-thin part $\tau_1$ with an exhausting sequence $(T_n)_{n}$ of $\mathbb{F}$-predictable stopping times.
\end{definition}
\vspace{.1em}
\begin{remark}\label{rem:increasing}
Let $\tau$ be a random time satisfying Hypothesis ($\bold{\mathcal{P}}$) w.r.t.~$\mathbb{F}$.~Then w.l.o.g.~the exhausting sequence $(T_n)_{n}$ can be chosen a.s.~increasing.
\end{remark}
\vspace{.1em}
\begin{proposition}\label{prop:0}
Any random time satisfies Hypothesis ($\bold{\mathcal{P}}$) w.r.t.~the natural Brownian filtration.
\end{proposition}
\begin{proof}
Any stopping time w.r.t.~the natural Brownian filtration is predictable.
\end{proof}
\vspace{.1em}
\begin{proposition}\label{prop:1}
Let $\tau$ be a random time satisfying Hypothesis ($\bold{\mathcal{P}}$) w.r.t.~$\mathbb{F}$.~Then:\\
(i)
$\tau_1$ is an $\mathbb{F}^{\tau_1}$-accessible stopping time with enveloping sequence $(T_n)_{n}$;\\
(ii)  $\tau_1$ and $\tau_2$ are the $\mathbb{F}^{\tau}$-accessible component and the  $\mathbb{F}^{\tau}$-totally inaccessible component of $\tau$, respectively, that is
$\tau_1=\tau^{a, \mathbb{F}^{\tau}}$ and $\tau_2=\tau^{i, \mathbb{F}^{\tau}}$.
\end{proposition}
\begin{proof}
(i) It follows immediately considering that any $\mathbb{F}$-predictable stopping time is also $\mathbb{F}^{\tau_1}$-predictable.\\
(ii) Previous point implies that $\tau_1$ is $\mathbb{F}^{\tau}$-accessible, since $\mathbb{F}^{\tau_1}\subset \mathbb{F}^{\tau}$.~Moreover, $\tau_2$ satisfies Hypothesis ($\bold{\mathcal{A}}$) w.r.t.~$\mathbb{F}^{\tau_1}$, since by definition $\tau_1$ and $\tau_2$ have disjoint graphs and $\tau_2$ avoids $\mathbb{F}$-stopping times.~By Remark \ref{rem:avoidance tot.in.ty}
 $\tau_2$ is $\big(\mathbb{F}^{\tau_1}\big)^{\tau_2}$-totally inaccessible, that is $\tau_2$ is $\mathbb{F}^{\tau}$-totally inaccessible (see Remark \ref{rem:composition}).
\end{proof}
\vspace{.1em}
\begin{remark}\label{rem:trivial}
Observe that $\tau$ satisfies Hypothesis ($\bold{\mathcal{P}}$) w.r.t.~$\mathbb{F}$ with $T_n=t_n\in\mathbb{R}^+, \; n\geq 1,$ if and only if $\tau_1$ is naturally accessible.~In fact $\tau_1$ is naturally accessible if and only if it is a discrete random variable (see Remark \ref{rem:predictability}) and any deterministic time is predictable w.r.t.~any filtration.\\
Moreover, when $\tau_1$ is naturally accessible, the (possibly finite) sequence of atoms of the law of $\tau_1$ is both a naturally enveloping sequence and an $\mathbb{F}$-exhausting sequence of $\tau_1$.
\end{remark}
\vspace{.1em}
\begin{proposition}\label{lemma:eq:Delta}
Assume Hypothesis ($\bold{\mathcal{P}}$) w.r.t.~$\mathbb{F}$ for $\tau$.Then
\begin{equation}\label{eq:delta}\{\Delta H^{\tau_1,\mathbb{F}^{\tau_1}}\neq 0\}\subset\bigcup_{n}\, [[T_n]].\end{equation}
%
%
\end{proposition}
\begin{proof}
Relation  (\ref{eq:delta}) derives immediately from the equality
\begin{equation}\label{eq-comp222}
H^{\tau_1, \mathbb{F}^{\tau_1}}_\cdot=\sum_{n}\big(\mathbb{I}_{C_n}-P(C_n\mid\mathcal{F}^{\tau_1}_{T_n^-})\big)\mathbb{
I}_{\{T_n\le
  \cdot\}}.
\end{equation}
The above representation  follows by Lemma 2.13 in \cite{caltor-accessible-23} joint with point (i) of Proposition \ref{prop:1} (see (\ref{def:C_n})).
\end{proof}
\section{An $\mathbb{F}^\tau$-basis}\label{sec3}
In this section, we prove the main theorem of this paper.~We work on a probability space $(\Omega,\mathbb{F},P)$ with a filtration $\mathbb{F}$ and a random time $\tau$.~Our theorem deals with constructing a basis on $\mathbb{F}^\tau$, that is on the progressive enlargement of $\mathbb{F}$ by a random time $\tau$.~The idea is to overcome the $\mathbb{F}$-avoidance
condition on $\tau$ allowing the $\mathbb{F}$-thin part of $\tau$ to be nontrivial and not an $\mathbb{F}$-stopping time.~The starting point is the equality $\sigma(\tau\wedge\cdot)=\sigma(\tau_1\wedge\cdot)\vee\sigma(\tau_2\wedge\cdot)$, which suggests looking at $\mathbb{F}^\tau$ as the progressive enlargement by $\tau_2$ of the filtration obtained progressively enlarging $\mathbb{F}$ by $\tau_1$.\bigskip\\
We consider the following setting.\vspace{.1em}\\
\textbf{$(\bold{\mathcal{S}})$}: \textit{$\mathbb{F}$ is a standard filtration with trivial initial $\sigma$-algebra;}
\textit{$(M^j)_{j\in J}$, $J\subset\mathbb{N}$, is an $\mathbb{F}$-basis (see Definition \ref{def:basis});}
\textit{$\tau$  is a random time which satisfies Hypothesis $(\bold{\mathcal{P}})$ w.r.t.~$\mathbb{F}$   (see Definition \ref{def:P}).}\vspace{.1em}\\
We need to introduce a \textit{continuity condition} on $\mathbb{F}$.\vspace{.1em}\\
Condition \textbf{$(\bold{\mathcal{C}})$}: \textit{For any $n\geq 1$, $\mathcal{F}_{T_n}=\mathcal{F}_{T_n^-}$.}
\vspace{.1em}
\begin{remark}
Condition \textbf{$(\bold{\mathcal{C}})$}  is equivalent to assuming that, for all $j\in J$ and for all $n\geq 1$, $\Delta M^j_{T_n}\equiv 0$.
\end{remark}
\vspace{.1em}
\begin{lemma}\label{cor}
    Let the setting \textbf{$(\bold{\mathcal{S}})$} be in force.~If $\mathbb{F}\hookrightarrow\mathbb{F}^\tau$, then all elements of the family $\big\{(M^j)_{j\in J}, H^{\tau_1,\mathbb{F}^{\tau_1}}\big\}$ are both $\mathbb{F}^{\tau_1}$-martingales and $\mathbb{F}^\tau$-martingales.~Moreover:\\
(i) \begin{equation}\label{eq02}\big[H^{\tau_1,\mathbb{F}^{\tau_1}}, H^{\tau_2,\mathbb{F}^\tau}\big]_\cdot\equiv 0\;\;\;a.s.\end{equation}
and therefore, $H^{\tau_1,\mathbb{F}^{\tau_1}}$ and  $H^{\tau_2,\mathbb{F}^{\tau}}$ are $\mathbb{F}^{\tau}$-orthogonal martingales;\\
(ii) if Condition \textbf{$({\mathcal{C}})$} holds, then, for all $j\in J$,
\begin{equation}\label{eq01}\big[H^{\tau_1,\mathbb{F}^{\tau_1}}, M^j\big]_\cdot\equiv 0\;\;\;a.s.\end{equation}
and therefore, $M^j$ and $H^{\tau_1,\mathbb{F}^{\tau_1}}$ are both $\mathbb{F}^{\tau_1}$-orthogonal martingales and $\mathbb{F}^{\tau}$-orthogonal martingales.
\end{lemma}
\begin{proof}
$\mathbb{F}\hookrightarrow\mathbb{F}^\tau$ implies both  $\mathbb{F}\hookrightarrow\mathbb{F}^{\tau_1}$ and $\mathbb{F}^{\tau_1}\hookrightarrow\mathbb{F}^\tau$, so that the $M^j$'s are both $\mathbb{F}^{\tau_1}$ and $\mathbb{F}^{\tau}$-martingales and $H^{\tau_1,\mathbb{F}^{\tau_1}}$ is an $\mathbb{F}^{\tau}$-martingale (see Proposition 5.4 in \cite{ak-chou-jean-18}).~By construction $H^{\tau_2,\mathbb{F}^\tau}$ is an $\mathbb{F}^\tau$-martingale.~Therefore, all covariation processes in (\ref{eq02}) and (\ref{eq01}) are well-defined.~Moreover, since $H^{\tau_1,\mathbb{F}^{\tau_1}}$ is a process of finite variation,
\begin{equation}\label{eq:cov1}\big[H^{\tau_1, \mathbb{F}^{\tau_1}}, H^{\tau_2,\mathbb{F}^\tau}\big]_\cdot=\sum_{s\leq \cdot}\Delta H^{\tau_1, \mathbb{F}^{\tau_1}}_s\,\Delta H^{\tau_2,\mathbb{F}^\tau}_s,\end{equation}
and
\begin{equation}\label{eq:cov2}\big[H^{\tau_1,\mathbb{F}^{\tau_1}},M^j\big]_\cdot=\sum_{s\leq \cdot}\Delta H^{\tau_1,\mathbb{F}^{\tau_1}}_s\,\Delta M^j_s.\end{equation}
(i) $H^{\tau_2,\mathbb{F}^\tau}$ jumps at $\tau_2$ (see Remark \ref{rem:avoidance tot.in.ty}) and Proposition \ref{lemma:eq:Delta} holds so that for all $t$
$$\Big\{\sum_{s\leq t}\Delta H^{\tau_1,\mathbb{F}^{\tau_1}}_s\,\Delta H^{\tau_2,\mathbb{F}^\tau}_s\neq 0\Big\}\;
\subset \;\bigcup_{n} \{\tau_2=T_n\}.$$
Therefore, (\ref{eq02}) follows by (\ref{eq:cov1})
and the $\mathbb{F}$-avoidance property of $\tau_2$.\bigskip\\
(ii) Proposition \ref{lemma:eq:Delta} states that $\{\Delta H^{\tau_1,\mathbb{F}^{\tau_1}}\neq 0\}$ is an $\mathbb{F}$-thin set (see Definition 3.18, p.89, in \cite{he-wang-yan92}) with exhausting sequence $(T_n)_{n\geq 1}$, while Condition \textbf{$(\mathcal{C})$} implies that any martingale $M^j$ cannot jump at any time $T_n$.~Hence, (\ref{eq01}) derives immediately from (\ref{eq:cov2}).
\end{proof}
\vspace{.1em}
\noindent We are now able to prove the main result of this paper.~In its proof, we will make use several times of Jacod-Yor's Lemma (see Theorem 7 in \cite{davis2}).\vspace{.1em}\\
\begin{thm}\label{thm:new}
Let the setting \textbf{$(\mathcal{S})$} and the Condition \textbf{$(\mathcal{C})$} be in force.\\
(i) If $\mathbb{F}\hookrightarrow\mathbb{F}^{\tau_1}$, then the family $\Big((M^j)_{j\in J},H^{\tau_1, \mathbb{F}^{\tau_1}}\Big)$ is an $\mathbb{F}^{\tau_1}$-basis.\\
(ii) If $\mathbb{F}\hookrightarrow\mathbb{F}^{\tau}$, then the family $\Big((M^j)_{j\in J}, H^{\tau,\mathbb{F}^\tau}\Big)$ is an $\mathbb{F}^\tau$-basis.
\end{thm}
\begin{proof}
  If $\tau_1$ is an $\mathbb{F}$-stopping time, then $\mathbb{F}=\mathbb{F}^{\tau_1}$ and $\mathbb{F^\tau}=\mathbb{F}^{\tau_2}$.~Therefore point (i) is trivial (see Remark \ref{rem pred}) and point (ii) is well-known (see Theorem 4.9 in \cite{ditella-eng-22}).\\
 Otherwise, we proceed as follows.\bigskip\\
(i) We sketch the line of the proof of this point.~We show that $P$ is the unique probability measure on $(\Omega, \mathcal{F}^{\tau_1}_\infty)$ that is a martingale measure for the family  $\Big((M^j)_{j\in J},H^{\tau_1, \mathbb{F}^{\tau_1}}\Big)$.~Then Jacod-Yor's Lemma implies that the stable subspace generated by the family $\Big((M^j)_{j\in J},H^{\tau_1, \mathbb{F}^{\tau_1}}\Big)$ coincides with $\mathcal{M}^2(P,\mathbb{F}^{\tau_1}_{\infty})$.~This ends the proof,
since the martingales of the family are pairwise orthogonal (see (\ref{eq01})).\bigskip\\
Let $Q$ be a probability measure on $(\Omega, \mathcal{F}^{\tau_1}_\infty)$ such that all elements of $\Big((M^j)_{j\in J},
H^{\tau_1, \mathbb{F}^{\tau_1}}\Big)$ belong to  $\mathcal{M}^2(Q, \mathbb{F}^{\tau_1})$.~By assumption $\mathcal{F}_0$ is trivial and  $(M^j)_{j\in J}$ is an $\mathbb{F}$-basis so that
Jacod-Yor's Lemma implies
\begin{equation}\label{eq:equiva1}
Q|_{\mathcal{F}_\infty}=P|_{\mathcal{F}_\infty}.
\end{equation}
Observe that $(M^j)_{j\in J}\in\mathcal{M}^2(Q, \mathbb{F}^{\tau_1})$ joint with $\mathbb{F}\hookrightarrow\mathbb{F}^{\tau_1}$ implies
\begin{equation}\label{eq:equiva2}
\mathbb{F}\hookrightarrow_{Q}\mathbb{F}^{\tau_1},
\end{equation}
that is the immersion property of $\mathbb{F}$ in $\mathbb{F}^{\tau_1}$ under $Q$.\bigskip\\
Since the uniqueness of the compensator of $\tau_1$, $H^{\tau_1, \mathbb{F}^{\tau_1}}_\cdot$ is an $\mathbb{F}^{\tau_1}$-martingale under $Q$ if and only if, for all $n\geq 1$,
\begin{equation}\label{eq:equa}
P(C_n\mid\mathcal{F}^{\tau_1}_{T_n^-})=Q(C_n\mid\mathcal{F}^{\tau_1}_{T_n^-})
\end{equation}
(see Lemma 13 in \cite{caltor-accessible-23} and (\ref{def:C_n})).\bigskip\\
The filtration $\mathbb{F}^{\tau_1}$ satisfies
\begin{equation}\label{eq:represen}
\mathcal{F}^{\tau_1}_t=\begin{cases}\mathcal{F}_t,  &  \text{ if } t<T_1\\
   \mathcal{F}_t\vee\sigma (C_1), & \text{ if } t\in [T_1,T_2)\\
   \vdots \\
   \mathcal{F}_t\vee\sigma (C_1,\ldots, C_n), & \text{ if } t\in [T_n,T_{n+1})\\
    \vdots
     \end{cases}
\end{equation}
and
\begin{equation}\label{eq:final}
\mathcal{F}^{\tau_1}_{\infty}=\bigvee_n \mathcal{F}_\infty\vee\sigma (C_1,\ldots, C_n)\end{equation}
(see Lemma 1.5  in \cite{ak-chou-jean-18}).\bigskip\\
The above recursive form of $\mathbb{F}^{\tau_1}$ suggests to apply an iterative procedure
to prove for any $n\geq 1$
\begin{equation}\label{eq:equiva2bis}
Q|_{\mathcal{F}_\infty\vee\sigma (C_1,\ldots, C_n)}=P|_{\mathcal{F}_\infty\vee \sigma (C_1,\ldots, C_n)}.
\end{equation}
The first step of the induction aims to show that
\begin{equation*}
Q|_{\mathcal{F}_\infty\vee\sigma (C_1)}=P|_{\mathcal{F}_\infty\vee \sigma (C_1)},
\end{equation*}
that is for any $A\in \mathcal{F}_\infty$
\begin{equation}\label{eq:step1}
P(A\cap C_1)=Q(A\cap C_1).
\end{equation}
Fixed $A\in \mathcal{F}_\infty$
\begin{equation*}
%
P(A\cap C_1)=E^P\big[P(A\cap C_1\mid \mathcal{F}_{T_1})\big]=E^Q\big[P(A\cap C_1\mid \mathcal{F}_{T_1})\big],
\end{equation*}
where the second equality follows by  (\ref{eq:equiva1}) using $\mathcal{F}_{T_1}\subset \mathcal{F}_{\infty}$.\bigskip\\
The immersion hypothesis $\mathbb{F}\hookrightarrow\mathbb{F}^{\tau_1}$ implies that, under $P$, $\mathcal{F}_{\infty}$ is conditionally independent of $\mathcal{F}^{\tau_1}_{T_1}$ given $\mathcal{F}_{T_1}$ (see Theorem 3 in \cite{Bre-yor78}).~Therefore, since $C_1\in \mathcal{F}^{\tau_1}_{T_1}$ (see (\ref{eq:represen}))
\begin{equation}\label{eq:equiva5}
P(A\cap C_1)=E^Q\big[P(A\cap C_1\mid \mathcal{F}_{T_1})\big]=E^Q\big[P(A\mid \mathcal{F}_{T_1})P(C_1\mid \mathcal{F}_{T_1})\big].
\end{equation}
By (\ref{eq:equiva1}), since $A\in \mathcal{F}_\infty$, it follows
\begin{equation}\label{eq: step11}
P(A\mid \mathcal{F}_{T_1})=Q(A\mid \mathcal{F}_{T_1}).\end{equation}
Condition \textbf{$(\mathcal{C})$} gives $\mathcal{F}_{T_1}=\mathcal{F}_{T_1^-}$ and therefore
\begin{equation}\label{eq: step12}P(C_1\mid \mathcal{F}_{T_1})=P(C_1\mid \mathcal{F}_{T_1^-})=Q(C_1\mid \mathcal{F}_{T_1^-})=Q(C_1\mid \mathcal{F}_{T_1}),\end{equation}
where the middle equality follows by (\ref{eq:equa}) with $n=1$.\bigskip\\
Using (\ref{eq: step11}) and (\ref{eq: step12}) equality (\ref{eq:equiva5}) becomes
\begin{equation}\label{eq:equiva6}
P(A\cap C_1)=E^Q\big[Q(A\mid \mathcal{F}_{T_1})Q(C_1\mid \mathcal{F}_{T_1})\big].
\end{equation}
Now, (\ref{eq:equiva2}) implies that the conditional independence of $\mathcal{F}^{\tau_1}_{T_1}$ given $\mathcal{F}_{T_1}$ also holds under $Q$, so that applying it in the right-hand side of (\ref{eq:equiva6}) one gets (\ref{eq:step1}).
\bigskip\\
The general step of the induction works similarly: fixed $n>1$, one assumes  %
\begin{equation*}
%
Q|_{\mathcal{F}_\infty\vee\sigma (C_1,\ldots, C_{n-1})}=P|_{\mathcal{F}_\infty\vee \sigma (C_1,\ldots, C_{n-1})}
\end{equation*}
and, for any choice of $A\in \mathcal{F}_\infty$ and $B\in \sigma (C_1,\ldots, C_{n-1})$, by the same technique one proves that $P(A\cap B\cap C_n)=Q(A\cap B\cap C_n)$.\\
Due to the arbitrariness of $A$ and $B$, the probability measures $P$ and $Q$ coincide on $\mathcal{F}_\infty\vee\sigma (C_1,\ldots, C_n)$, that is (\ref{eq:equiva2bis}).\bigskip\\
Finally, since (\ref{eq:final}), a monotone class argument yields that  $P$ and $Q$ coincide on $\mathcal{F}^{\tau_1}_{\infty}$.\bigskip\\
(ii) $\mathbb{F}\hookrightarrow\mathbb{F}^\tau$ implies $\mathbb{F}\hookrightarrow\mathbb{F}^{\tau_1}$ and previous point holds.\\
The random time $\tau_2$ satisfies Hypothesis ($\mathcal{A}$) w.r.t.~$\mathbb{F}^{\tau_1}$.~In fact $\tau_2$ by definition avoids both  $\tau_1$, and all $\mathbb{F}$-stopping times.~Moreover, $\mathbb{F}^{\tau_1}\hookrightarrow\mathbb{F}^\tau$ (see Proposition 5.4 in \cite{ak-chou-jean-18}).\\
Since $\big(\mathbb{F}^{\tau_1}\big)^{\tau_2}=\mathbb{F}^{\tau}$ (see Remark \ref{rem:composition}), we can apply  Theorem 4.9 of \cite{ditella-eng-22} to the progressive enlargement of $\mathbb{F}^{\tau_1}$ by $\tau_2$.~It follows that  $$\Big((M^j)_{j\in J}, H^{\tau_1, \mathbb{F}^{\tau_1}}, H^{\tau_2,\mathbb{F}^\tau}\Big)$$ is an $\mathbb{F}^\tau$-basis.\bigskip\\
It is to note that $\mathbb{F}^{\tau_1}\hookrightarrow\mathbb{F}^\tau$ implies
\begin{equation}\label{eq-comp223}
H^{\tau_1,\mathbb{F}^{\tau_1}} = H^{\tau_1,\mathbb{F}^\tau}.\end{equation}
 Thus, we have proved that $\Big((M^j)_{j\in J}, H^{\tau_1, \mathbb{F}^{\tau}}, H^{\tau_2,\mathbb{F}^\tau}\Big)$ is an $\mathbb{F}^\tau$-basis.~Therefore for any $V\in \mathcal{M}^2(P,\mathbb{F}^\tau)$
\begin{align}\label{eq:basis2}
V_t=V_0+\sum_{i}\int_0^t \Phi^V_j(s) d M^j_s+\int_0^t \gamma(s) d H^{\tau_1,\mathbb{F}^\tau}_s+\int_0^t \eta(s) d H^{\tau_2,\mathbb{F}^\tau}_s,\;\;t\geq 0.
\end{align}
Here, $V_0$ is a random variable $\mathcal{F}_0$-measurable,
$\Phi^V_j$, for all $j\in J$, $\gamma$ and $\eta$ are $\mathbb{F}^\tau$-predictable processes such that the random variables $$\int_0^\infty\,(\Phi^V_j)^2(s) d[M^j]_s,\;\;\int_0^\infty\,\eta^2(s) d[ H^{\tau_1,\mathbb{F}^\tau}]_s,\;\;\int_0^\infty\,\gamma^2(s) d[H^{\tau_2,\mathbb{F}^\tau}]_s$$ have finite expectations (see Definition \ref{def:basis}).\bigskip\\
Let $\mathcal{D}$ be the  $\mathbb{F}^\tau$-predictable subset of $\Omega \times [0,T]$ defined as
$$\mathcal{D}:=\bigcup_n\,[[T_n]].$$
Now,
 \begin{equation}\label{eq:rangle}\int_{\overline{\mathcal{D}}}d\langle H^{\tau_1,\mathbb{F}^\tau}\rangle_t =0\;\;\;\;\;\; \int_{\mathcal{D}}d\langle H^{\tau_2,\mathbb{F}^\tau}\rangle_t =0.\end{equation}
The first equality above follows by Proposition 2.15 in \cite{caltor-accessible-23}, (\ref{eq-comp223}) and (\ref{eq-comp222}).~The second one, instead, follows observing
that  $\langle H^{\tau_2,\mathbb{F}^\tau}\rangle$ is a continuous process, since $\tau_2$ is $\mathbb{F}^\tau$-totally inaccessible (see Proposition \ref{prop:1} (ii) and Theorem 7.11, p.195, in \cite{he-wang-yan92}).\bigskip\\
 Let $\rho$ be the predictable process defined by
 $$\rho_t:=\gamma_t \mathbb{I}_{\mathcal{D}}(t)+\eta_t \mathbb{I}_{\overline{\mathcal{D}}}(t).$$
 It holds $E\left[\int_0^\infty\,\rho^2_s d[ H^{\tau,\mathbb{F}^\tau}]_s\right]<+\infty$.\bigskip\\
  By (\ref{eq:rangle}), $\mathcal{D}$ is a predictable support of $d\langle H^{\tau_1,\mathbb{F}^\tau}\rangle$ and $\overline{\mathcal{D}}$ is a predictable support of $d\langle H^{\tau_2,\mathbb{F}^\tau}\rangle$.~Following the same lines as in the proof of Proposition 4.2 (i) in \cite{caltor-accessible-23}, we derive
  \begin{equation*}\int_0^t \gamma_s d H^{\tau_1,\mathbb{F}^\tau}_s=\int_0^t \rho_s d H^{\tau_1,\mathbb{F}^\tau}_s,\;\;\;\;\;   \int_0^t \eta_s d H^{\tau_2,\mathbb{F}^\tau}_s=\int_0^t \rho_s d H^{\tau_2,\mathbb{F}^\tau}_s,\;t\geq 0.\end{equation*}
  Therefore,
since
 \begin{equation}\label{eq:sumH}
 H^{\tau,\mathbb{F}^\tau}= H^{\tau_1,\mathbb{F}^\tau} + H^{\tau_2,\mathbb{F}^\tau},
 \end{equation}
equality (\ref{eq:basis2}) can be rewritten as
  \begin{align*}
V_t=V_0+\sum_{i}\int_0^t \Phi^V_j(s) d M^j_s+\int_0^t \rho(s) d H^{\tau,\mathbb{F}^\tau}_s,\;\;t\geq 0.
\end{align*}
 By the arbitrariness of $V$ the last equality gives the thesis.
\end{proof}
\vspace{0.1em}
\begin{remark}\label{rem:generalized}
Let us recall that whatever $\tau$ is, it holds
$H^{\tau_2,\mathbb{F}^\tau}=\mathbb{I}_{\tau_2\leq\cdot}-\Lambda^{\tau_2\wedge\cdot},$ where $\Lambda$ denotes the \textit{$\mathbb{F}^{\tau_1}$-reduction of the compensator of $\tau_2$} (see e.g.~Proposition 2.11, p.36, in \cite{aksamit-jean-17}).~Thus, in the setting of point (ii) of the above theorem,  using (\ref{eq:sumH}), (\ref{eq-comp223}) and (\ref{eq-comp222}) one gets
\begin{equation}\label{eq:compen}
H^{\tau,\mathbb{F}^\tau}=\mathbb{I}_{\tau\leq\cdot}-\sum_{n}P(C_n\mid\mathcal{F}^{\tau_1}_{T_n^-})\mathbb{
I}_{\{T_n\le
  \cdot\}}-\Lambda^{\tau_2\wedge\cdot}.
\end{equation}
In particular, when $\tau_2$ satisfies \textit{Jacod's density hypothesis} w.r.t.~$\mathbb{F}^{\tau_1}$ with \textit{intensity $\lambda$} (see Condition (A) and Proposition 1.5 in \cite{jacod85}),
\begin{equation}\label{eq:generalized}
H^{\tau,\mathbb{F}^\tau}=\mathbb{I}_{\tau\leq\cdot}-\sum_{n}P(C_n\mid\mathcal{F}^{\tau_1}_{T_n^-})\mathbb{
I}_{\{T_n\le
  \cdot\}}-\int^{\tau_2\wedge\cdot}_0\lambda_s\;ds.
\end{equation}
Therefore, when the exhausting sequence of $\tau_1$ is finite, $\tau$ is under the \textit{generalized density hypothesis} (see  \cite{jiao-li15}).
\end{remark}
\section{Examples}\label{sec4}
In the literature, many models dealing with progressive enlargement take as reference filtration the natural filtration of a L\'evy process.~For the sake of clarity, we restate Theorem \ref{thm:new} in this particular case.~It is worth noting that the natural filtration of any L\'evy process always admits a basis (see Theorem 4.3 in  \cite{ditella-eng-22}).~Let us recall that the process $H^{\tau,\mathbb{F}^\tau}$ under the hypotheses of point (ii) of the theorem has the form (\ref{eq:compen}).\vspace{.1em}\\
\begin{thm}\label{cor:Levy}
 Let $\mathbb{F}$ be the natural filtration of a L\'evy process and $(M^j)_{j\in J}, J\subset\mathbb{N},$ any $\mathbb{F}$-basis.\\
 Let $\tau$ be any random time that satisfies Hypothesis $(\bold{\mathcal{P}})$ and let $H^{\tau,\mathbb{F}^\tau}$ be the $\mathbb{F}^\tau$-compensated occurrence process of $\tau$.\\
 If $\mathbb{F}\hookrightarrow\mathbb{F}^\tau$, then $\Big((M^j)_{j\in J}, H^{\tau,\mathbb{F}^\tau}\Big)$ is an $\mathbb{F}^\tau$-basis.
\end{thm}
\begin{proof}
$\mathbb{F}$ is quasi-left continuous  and therefore Condition \textbf{$(\mathcal{C})$} holds (see Proposition 7, p.21, in \cite{MR1406564}).~Thus, Theorem \ref{thm:new} applies.
\end{proof}
\vspace{.1em}
\noindent As an immediate consequence of the previous theorem and Proposition \ref{prop:0}, we get the next result.~It deals with the natural generalization of the martingales representation property of the Brownian filtration.\vspace{.1em}\\
\begin{cor}\label{cor:Brow}
 Let $\mathbb{F}$ be the natural filtration of a Brownian motion $W$ and $\tau$ any random time.~If  $\mathbb{F}\hookrightarrow\mathbb{F}^\tau$ holds, then $\Big(W,H^{\tau,\mathbb{F}^\tau}\Big)$ is an $\mathbb{F}^\tau$-basis.
\end{cor}
\vspace{.1em}
\begin{remark}
\noindent The \textit{hybrid sovereign default model} proposed  by \cite{MR3758923} falls among the possible examples of application of the above corollary.~In that model the default is defined as
$$\tau:=\zeta^*\wedge\xi.$$
The random time $\zeta^*$ describes \textit{successive solving downgrades} due to economic and political inferences, while $\xi$ stays for the \textit{idiosyncratic risks}.~From a mathematical point of view, $\zeta^*$ is an accessible random time with a finite number of predictable components and  $\xi$  is a totally inaccessible random time.~The process driving the \textit{solvency process} is a Brownian motion,  $\tau_1$ is equal to $\zeta^*$ and $\tau_2$ is equal to $\xi$.~The authors show that the immersion property holds.~Therefore, Corollary \ref{cor:Brow} applies.
\end{remark}
\vspace{0.5em}
%
%
%
%
\noindent The generalization of the classical Cox procedure for default times  provides a class of applications of  Theorem \ref{cor:Levy} (see \cite{MR4529908}).\vspace{.1em}\\
\begin{cor}
Let $\mathbb{F}$ be the natural filtration of a L\'evy process and $(M^j)_{j\in J}, J\subset\mathbb{N},$ any $\mathbb{F}$-basis.~Let $\tau$ be defined by
$$\tau:=\inf\{t\geq 0, K_t\geq \Theta\},$$
where $K$ is a c\`adl\`ag $\mathbb{F}$-predictable process such that $K_0=0$, and $\Theta$ is a random variable  with exponential law of parameter one, independent of  $\mathcal{F}_\infty$.\\Then $\Big((M^j)_{j\in J}, H^{\tau,\mathbb{F}^\tau}\Big)$ is an  $\mathbb{F}^\tau$-basis.
\end{cor}
\begin{proof}
  Theorem \ref{cor:Levy} applies.~In fact: the immersion property of $\mathbb{F}$ in $\mathbb{F}^\tau$ is obtained for free by Cox construction (see e.g.~p.471 in \cite{MR4529908}); the  Hypothesis $(\bold{\mathcal{P}})$ for $\tau$ derives from the assumptions on $K$ (see Section 3.2.5, p.477, in \cite{MR4529908}).
\end{proof}
\vspace{.1em}
\begin{remark}
 If the continuous part of the process $K$ in the above corollary is  equal to $\int_0^\cdot \lambda_s ds$, with $\lambda$ an $\mathbb{F}$-adapted  positive process, then $H^{\tau,\mathbb{F}^\tau}$ has the expression (\ref{eq:generalized}).
\end{remark}
\vspace{.1em}
\noindent In the following model, dealing with a special class of Brownian filtrations, the random time is not constructed by Cox procedure.~This example has been proposed in \cite{MR4152276} studying the market's completeness under filtration shrinkage and in \cite{ak-chou-jean-18} as an example of thin time.\bigskip\\
Let $B$ be the \textit{L\'evy transformation of a Brownian motion} $W$, that is the process $$B:=\int_0^\cdot sign(W_s)\;dW_s$$ and let $\mathbb{F}^B$ and $\mathbb{F}^W$ be the natural filtrations of $B$ and $W$, respectively.~As well-known  $B$ is a natural Brownian motion,  and $\mathbb{F}^B=\mathbb{F}^{|W|}\subsetneq\mathbb{F}^W$, and the exponential martingale  $$S:=\mathcal{E}(B)$$ is both an $\mathbb{F}^B$-basis and an $\mathbb{F}^W$-basis.~In the language of finance modeling, $(S,\mathbb{F}^B)$ is a complete market, as well as $(S,\mathbb{F}^W)$.~But, when
\begin{equation}\label{def:tau ex}\tau:=\inf\{t\geq 0: W_t=1\},\end{equation}
$S$ is not an $(\mathbb{F}^B)^\tau$-basis.\\
In fact, if $(T_n)_{n\geq 1}$ is the sequence of $\mathbb{F}^B$-stopping times defined by
$$T_n:=\inf\{t> S_{n-1}: |W_t|=1\}$$
with
$$S_n:=\inf\{t> T_{n-1}: |W_t|=0\},\;n>1,\;\;\;\text{and}\;\;\;S_0:=0,$$
then the process
\begin{equation*}
N:=\mathbb{
I}_{\{\tau\le
  \cdot\}}-\frac12\sum_{n\geq 1}\mathbb{I}_{\tau\geq T_n}\mathbb{
I}_{\{T_n\le
  \cdot\}},
\end{equation*}
is a discontinuous $(\mathbb{F}^B)^\tau$-martingale and cannot be represented by integration w.r.t.~$S$.\bigskip\\
In \cite{MR4152276}, since $\mathbb{F}^B\subset \big(\mathbb{F}^B\big)^\tau\subset\mathbb{F}^W$, this example is useful to explain why some markets could lose completeness after the reduction of information or, vice versa, when the addition of information is too little.\\
In \cite{ak-chou-jean-18}, by the previous example, the authors discuss a case of a natural totally inaccessible random time, $\tau$, which coincides with its thin part w.r.t.~a given filtration, $\mathbb{F}^B$, with an exhausting sequence of predictable stopping times, $(T_n)_{n\geq 1}$.~We stress that $\tau$ is $(\mathbb{F}^B)^\tau$-accessible and becomes  predictable in the even larger filtration $\mathbb{F}^W$.\bigskip\\
Here we answer the question:  how can be represented all $(\mathbb{F}^B)^\tau$-local martingales since $S$ is not enough?\vspace{.1em}\\
\begin{proposition} $(S,N)$ is an $(\mathbb{F}^B)^\tau$-basis.\end{proposition}
\begin{proof}
We apply Theorem \ref{thm:new} to $\mathbb{F}^B$, $S$ as $\mathbb{F}^B$-basis and $\tau$ given by (\ref{def:tau ex}), and, more precisely, since $\tau$ is an $\mathbb{F}^B$-thin time, we apply point (i) of Theorem \ref{thm:new}.\bigskip\\
$\mathbb{F}^B\hookrightarrow (\mathbb{F}^B)^\tau$, since $\mathbb{F}^B\hookrightarrow\mathbb{F}^W$.\\
$\mathbb{F}^B$, like any Brownian natural filtration, satisfies Condition \textbf{$(\mathcal{C})$}.\\
It remains to check that $H^{\tau, (\mathbb{F}^B)^{\tau}}$, for which the analogous of (\ref{eq-comp222}) holds, coincides with $N$.\bigskip\\
Let us observe that
$$P(C_n\mid (\mathcal{F}^B)^\tau_{T_n^-})=P(C_n\cap\{\tau\geq T_n\}\mid (\mathcal{F}^B)^\tau_{T_n^-})=P(C_n\mid (\mathcal{F}^B)^\tau_{T_n^-})\mathbb{I}_{\tau\geq T_n},$$
where last equality follows by $$\{\tau< T_n\} \in (\mathcal{F}^B)^\tau_{T_n^-}.$$
The analogous of (\ref{eq-comp222}) reads
\begin{equation*}
H^{\tau, (\mathbb{F}^B)^\tau}_\cdot=\mathbb{
I}_{\{\tau\le
  \cdot\}}-\sum_{n\geq 1}P\left(C_n\mid(\mathcal{F}^B)^\tau_{T_n^-}\right)\mathbb{I}_{\tau\geq T_n}\mathbb{
I}_{\{T_n\le
  \cdot\}}.
\end{equation*}
In fact, for any fixed $n\geq 1$,
\begin{equation*}
P\big(C_n\mid(\mathcal{F}^B)^\tau_{T_n^-})\mathbb{I}_{\tau\geq T_n}=P(W_{T_n}=1\mid\mathcal{F}^{|W|}_{T_n}\vee\sigma (C_1,\ldots, C_{n-1})\big)\mathbb{I}_{\tau\geq T_n},\end{equation*}
where we have used (\ref{eq:represen}) and $$\mathcal{F}^{|W|}_{T_n^-}=\mathcal{F}^{B}_{T_n^-}=\mathcal{F}^{B}_{T_n}=\mathcal{F}^{|W|}_{T_n}.$$
Moreover,
$$P\big(W_{T_n}=1\mid\mathcal{F}^{|W|}_{T_n}\vee\sigma (C_1,\ldots, C_{n-1})\big)\mathbb{I}_{\tau\geq T_n}=\frac12\;\mathbb{I}_{\tau\geq T_n}.$$
To prove last equality, we observe that
$$P\big((W_{T_n}=1)^c\mid\mathcal{F}^{|W|}_{T_n}\vee\sigma (C_1,\ldots, C_{n-1})\big)=P\big(-W_{T_n}=1\mid\mathcal{F}^{|-W|}_{T_n}\vee\sigma (C_1,\ldots, C_{n-1})\big)$$
and, setting $D_h:=\{W_{T_h}=1\}=\{-W_{T_h}=1\}^c,\;h=1,\ldots,n-1,$ that
$$\sigma (C_1,\ldots, C_{n-1})=\sigma (D_1,\ldots, D_{n-1}).$$
Finally, by the symmetry of Brownian motion,
$$P\big(W_{T_n}=1\mid\mathcal{F}^{|W|}_{T_n}\vee\sigma (C_1,\ldots, C_{n-1})\big)=P\big((W_{T_n}=1)^c\mid\mathcal{F}^{|W|}_{T_n}\vee\sigma (C_1,\ldots, C_{n-1})\big).$$
\end{proof}
\section{Comments and perspectives}
The proof of our main theorem relies on two facts: the recursive formula of the filtration obtained progressively enlarging the reference filtration by a thin time (see (\ref{eq:represen})) and the expression of the compensated occurrence process of any accessible stopping time (see (\ref{eq-comp222})).\\

The object of ongoing research is, using the same ideas, to derive martingale representations relaxing the assumption on the thin part of $\tau$ and in the framework of multiple defaults.

\bibliography{biblio}

\end{document}